An object is placed first at infinity and then at 20 cm from the object side focal plane of a convex lens the two images thus form are 5 CM apart the focal length of the lens is

Grade:9

1 Answers

Arun
6 years ago
Dear Pranay
For first case
1/f = 1/v - 1/infinity
f = v
For second case
1/f = 1/(f+5)  -  1/-(f+20)
On solving
f = 10 cm
Regards
Arun (askIITians forum expert)